OpenStudy (anonymous):

Finding the compound function f'(f^-1(x))?

OpenStudy (anonymous):

\[The function is f(x)=\sqrt{x^2+3}, x \epsilon [-2,-1]\]

OpenStudy (anonymous):

and im wanting to find\[f'(f^-1(x))\]

zepdrix (zepdrix):

Hmmmmm ok let's start with the inner function and work our way outward.

OpenStudy (anonymous):

ok thats what i did, must of made an error

zepdrix (zepdrix):

\[\Large f(x)=\sqrt{x^2+3}\]\[\Large f^{-1}(x)=?\]

OpenStudy (anonymous):

i got -sqrt(x^2-3) for f^-1(x)

OpenStudy (anonymous):

then i wanted to differentiate that

zepdrix (zepdrix):

Mmmmmmmmm yah I think you're on the right track so far.

OpenStudy (anonymous):

so i differentiated and got \[\frac{ -x }{ \sqrt{x^2-3} }\]

zepdrix (zepdrix):

And that was wrong? :o

OpenStudy (anonymous):

supposedly its not right

zepdrix (zepdrix):

Try it without the minus sign. If that's correct, I can try to explain why it is such :d

OpenStudy (anonymous):

its not right, cause for \[\frac{ d }{ dx }f^-1(x)= \frac{ -x }{ \sqrt{x^2-3} }\]

OpenStudy (anonymous):

but im wanting to find \[f'(f^-1(x))\]

OpenStudy (anonymous):

the compound function

zepdrix (zepdrix):

Oh oh oh i see :D

zepdrix (zepdrix):

Oh ok ok ok ok. So here's what we'll do...

zepdrix (zepdrix):

We found the inner portion, good good good. Our next step, find f'(x). AFTERRRR we do that, we can evaluate the derivative function at f^{-1}(x)

OpenStudy (anonymous):

are we putting f^-1(x) into the derivative of f(x)

zepdrix (zepdrix):

yes, but only AFTERRRRRR you've taken the derivative of f(x) :)

OpenStudy (anonymous):

ok so derivative of f(x)=x/sqrt(x^2+3)

zepdrix (zepdrix):

So I guess we didn't need to take the derivative of the inverse :P that was silly hehe

zepdrix (zepdrix):

kk

zepdrix (zepdrix):

\[\Large f'(\color{orangered}{f^{-1}(x)})=\frac{\color{orangered}{f^{-1}(x)}}{\sqrt{\color{orangered}{f^{-1}(x)}^2+3}}\]

zepdrix (zepdrix):

A bit tricky to plug all that in, hmm

OpenStudy (anonymous):

ok and then simplify?

zepdrix (zepdrix):

Hmm ya i guess so +_+

zepdrix (zepdrix):

\[\Large f'(\color{orangered}{f^{-1}(x)})\quad=\quad \frac{\color{orangered}{\dfrac{-x}{\sqrt{x^2-3}}}}{\sqrt{\left(\color{orangered}{\dfrac{-x}{\sqrt{x^2-3}}}\right)^2}+3}\]

zepdrix (zepdrix):

woops i didn't put the 3 under the root -_- oh well lol work it out! :D

OpenStudy (anonymous):

ok thanks i understand this now